# Adaptively Robust and Sparse

### v1

@@ Abstract
While $K$-means is known to be a standard clustering algorithm, it may be compromised due to the presence of outliers and high-dimensional noisy variables. This paper proposes adaptively robust and sparse $K$-means clustering (ARSK) to address these practical limitations of the standard $K$-means algorithm. We introduce a redundant error component for each observation for robustness, and this additional parameter is penalized using a group sparse penalty. To accommodate the impact of high-dimensional noisy variables, the objective function is modified by incorporating weights and implementing a penalty to control the sparsity of the weight vector. The tuning parameters to control the robustness and sparsity are selected by $\rm Gap$ statistics. Through simulation experiments and real data analysis, we demonstrate the superiority of the proposed method to existing algorithms in identifying clusters without outliers and informative variables simultaneously.
Key words: Clustering; $K$-means; Variable selection; Gap statistic; Robust statistics
@@ Introduction
Identifying clustering structures is recognized as an important task in pursuing potential heterogeneity in datasets. While there are a variety of clustering methods, $K$-means clustering (Forgy, 1965) is the most standard clustering algorithm and is widely used in many scientific applications. However, real-world datasets present significant difficulties, such as the presence of outliers and high-dimensional noisy characteristics. Regarding the outlier issues, there are some attempts to robustify the standard $K$-means such as trimmed $K$-means (Cuesta-Albertos et al., 1997) and robust $K$-means (Klochkov et al., 2021). On the other hand, existence of high-dimensional noisy characteristics is typically addressed through clustering approaches that focus on sparsity for variable selection. In particular, Witten and Tibshirani (2010) proposed a lasso-type penalty to the variable weights incorporated in the clustering objective function. However, the efficacy of this algorithm might be lost if the dataset contains a significant number of outliers. While there are several methods to address either of two aspects (existence of outliers and noisy variables), practically useful methods to address two aspects simultaneously are scarce. One exception is the method proposed by Kondo et al. (2016). However, this approach requires the assumption of a trimming level in advance. In contrast, Brodinová et al. (2019) Introduces an approach beginning with eliminating outliers identified by the model. However, some helpful information can get thrown out with the outliers. Moreover, the model learns to handle unusual data better by adjusting outliers instead of removing them. This helps make it more robust with new data. This is particularly crucial when dealing with real-world data.
@@
Apart from clustering methods based on objective functions, probabilistic clustering using mixture models is also popular. It is typically done by fitting multivariate Gaussian mixtures. While several approaches have been proposed for stable estimation and clustering under existence of outliers (e.g. Coretto and Hennig, 2016; Fujisawa and Eguchi, 2006; Punzo and McNicholas, 2016; Sugasawa and Kobayashi, 2022; Yang et al., 2012), such approach suffers from a huge number of parameters (especially for modeling covariance matrix) to be estimated when the dimension is large. Therefore, these methods would not be a reasonable solution when our primary focus is on clustering.
@@
To solve the difficulty of a series of sparse trimming-cluster and probabilistic-cluster approaches, we propose a new approach based on regularization techniques for sparsity as well as robustness, called adaptively robust and sparse $K$-means (ARSK). We introduce an error component for each observation as an additional parameter to absorb the undesirable effects caused by outliers, making the clustering algorithm robust. We employ a group lasso penalty and a group SCAD penalty to estimate these parameters during clustering steps. Meanwhile, to reduce the data’s dimension, we also introduce a lasso-type penalty or SCAD penalty to weights for each variable to introduce sparsity and eliminate noise variables irrelevant to clustering. We then develop an efficient computation algorithm to minimize the proposed objective function. While the proposed method has two tuning parameters controlling robustness (the number of outliers to be detected) and sparsity (the level of the sparsity of the variable weight), we adopt a modified version of the $\rm Gap$ statistics (Tibshirani et al., 2001) to determine the optimal values of them. Hence, the level of robustness and sparsity is adaptively determined in a fully data-dependent manner.
@@
The paper is organized as follows. Section 2 reviews the $K$-means algorithm and its extended model introduces the proposed method ARSK and presents its details and pseudocode. Section 3 also presents the selection approach of tuning parameters and its pseudocode. Section 4 discusses ARSK’s robustness, while Section 5 presents the tuning parameter selection result. The two algorithms are applied to various scenarios of artificial datasets and several real-world datasets. This paper ends in Section 5 with concluding remarks. R code implementing the proposed method is available at GitHub repository (https://github.com/lee1995hao/ARSK).
@@
Suppose that we observe a $p$-dimensional vector, $x_{i\cdot}=(x_{i1},\ldots,x_{ip})$, for $i=1,\ldots,n$, and that we are interested in clustering $x_{i\cdot}$. Given the number of clusters, $K$, the conventional $K$-means algorithm is to find the optimal clustering by minimizing the following objective function:
$\min_{c_{1},...,c_{K}}\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{{i},{i^{\prime}}\in c_{k}}(x_{i\cdot}-x_{i^{\prime}\cdot})^{2},$ (1)
where $\|x_{i\cdot}-x_{i^{\prime}\cdot}\|$ is the $L_{2}$-distance between two observed vectors. Here $c_{k}$ is a set of indices representing a cluster of observations and $n_{k}$ is the size of $c_{k}$, such that $\cup_{k=1}^{K}c_{k}=\{1,\ldots,n\}$ and $\sum_{k=1}^{K}n_{k}=n$, where $c_{k}\cap c_{k^{\prime}}=\phi$ for $k\neq k^{\prime}$. The objective function (1) can be easily optimized by iteratively updating cluster means and assignment (e.g. Lloyd, 1982).
As demonstrated in Witten and Tibshirani (2010), the objective function (1) can be reformulated in terms of the between-cluster sum of squares, given by $\max_{c_{1},..,c_{k}}\sum_{j=1}^{p}Q_{j}(x_{\cdot j};\Theta)$, where
$Q_{j}(x_{\cdot j};\Theta)=\frac{1}{n}\sum_{i=1}^{n}\sum_{i^{\prime}=1}^{n}(x_{ij}-x_{i^{\prime}j})^{2}-\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{i,i^{\prime}\in c_{k}}(x_{ij}-x_{i^{\prime}j})^{2},$ (2)
with $x_{\cdot j}=(x_{1j},\ldots,x_{nj})$ and $\Theta$ being a collection of $c_{1},\ldots,c_{K}$. Witten and Tibshirani (2010) employed the modified version of the objective function (2), given by $\sum_{j=1}^{p}\omega_{j}Q_{j}(x_{\cdot j};\Theta)$ with an appropriate constrained on $\omega_{j}$, for conducting variable selection and clustering simultaneously.
@@
The objective function (2) is sensitive to outliers since the approach relies on the $L_{2}$ distance between each observation. As a result, even a small quantity of deviating observations can affect the $K$-means algorithm. In order to robustify the $K$-means approach, we modify the function $Q_{j}(x_{\cdot j};\Theta)$ as follows:
$\centering\begin{split}Q_{j}^{R}(x_{\cdot j};\Theta,E)&=\frac{1}{n}\sum_{i=1}^{n}\sum_{i^{\prime}=1}^{n}\{(x_{ij}-E_{ij})-(x_{i^{\prime}j}-E_{i^{\prime}j})\}^{2}\\ &\ \ \ -\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{i,i^{\prime}\in c_{k}}\{(x_{ij}-E_{ij})-(x_{i^{\prime}j}-E_{i^{\prime}j})\}^{2},\end{split}\@add@centering$ (3)
where $E$ is an $n\times p$ error matrix, which is a collection of $E_{ij}$ and $E_{ij}$ is an additional location parameter for each $x_{ij}$ such that $E_{ij}$ will be non-zero values for outlying observations $x_{ij}$ to prevent an undesirable effects from outliers. While the number of elements of $E$ is the same as the number of observations, it can be assumed that $E$ is sparse in the sense that most elements of $E$ are 0. That is, most elements are not outliers. To stably estimate $E$, we consider penalized estimation for $E_{ij}$ when optimizing the function (3). Such approaches are used in the robust fitting of regression models (e.g. She and Owen, 2011; Katayama and Fujisawa, 2017).
@@
$L(\Theta,E,W;\lambda)=\sum_{j=1}^{p}w_{j}Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j})-\sum_{i=1}^{n}P_{1}(\|E_{i\cdot}\|_{2};\lambda_{1})-\sum_{j=1}^{p}(P_{2}(w_{j};\lambda_{2})+\frac{1}{2}w_{j}^{2}),$ (4)
where $P_{1}$ and $P_{2}$ are group penalty functions and penalty functions for sparse (shrinkage) estimation of $E_{i\cdot}$ and $w_{j}$, respectively, and $\lambda=(\lambda_{1},\lambda_{2})$ is a set of tuning parameters. To find a unique solution for the weight coefficients $w_{j}$, we add a quadratic term $\frac{1}{2}\sum_{j=1}^{p}w_{j}^{2}$ to the loss function. Then, the robust and sparse $K$-means clustering can be defined as $\widehat{\Theta}$ such that
$\begin{split}(\widehat{\Theta},\widehat{E},\widehat{W})={\rm argmax}_{\Theta,E,W\in\mathcal{H}}L(\Theta,E,W;\lambda)\\ \mathcal{H}=\left\{{w_{j}\in\mathbb{R}^{+}:\sqrt{\sum_{j=1}^{p}w^{2}_{j}}=1}\right\}.\end{split}$
In what follows, we employ two specific forms of group penalty functions $P_{1}(\|E_{i\cdot}\|_{2};\lambda_{1})$: one type is the convex group penalty function, which is the group lasso penalty as $\lambda_{1}\|E_{i\cdot}\|_{2}$ (Yuan and Lin, 2006), and the other type is the nonconvex group penalty function, which is the group SCAD penalty (Huang et al., 2012) as
$P_{\lambda_{1}}^{SCAD}\left(\|E_{i\cdot}\|_{2}\right)=\left\{\begin{array}[]{ll}\lambda_{1}\|E_{i\cdot}\|_{2}&\text{ if }\|E_{i\cdot}\|_{2}\leq\lambda_{1};\\ -\left(\frac{\|E_{i\cdot}\|_{2}^{2}-2a\lambda_{1}\|E_{i\cdot}\|_{2}+\lambda_{1}^{2}}{2(a-1)}\right)&\text{ if }\lambda_{1}<\|E_{i\cdot}\|_{2}\leq a\lambda_{1};\\ \frac{(a+1)\lambda_{1}^{2}}{2}&\text{ if }\|E_{i\cdot}\|_{2}>a\lambda_{1}.\end{array}\right.$
@@
$P_{\lambda_{2}}^{SCAD}\left(w_{j}\right)=\left\{\begin{array}[]{ll}\lambda_{2}\left|w_{j}\right|&\text{ if }\left|w_{j}\right|\leq\lambda_{2};\\ -\left(\frac{\left|w_{j}\right|^{2}-2a\lambda_{2}\left|w_{j}\right|+\lambda_{2}^{2}}{2(a-1)}\right)&\text{ if }\lambda_{2}<\left|w_{j}\right|\leq a\lambda_{2};\\ \frac{(a+1)\lambda_{2}^{2}}{2}&\text{ if }\left|w_{j}\right|>a\lambda_{2}.\end{array}\right.$
When applying the $L_{1}$ or SCAD penalties, we seek the optimal weight vector of $W=(w_{1},\ldots,w_{p})$, which is normalized to have a unit L2-norm, i.e., $\sqrt{\sum_{j=1}^{p}w^{2}_{j}}=1$. This approach helps to promote a more balanced solution (Zou and Hastie, 2010).
@@
The proposed objective function (4) can be optimized by iteratively updating each parameter. Regarding the updating of the clustering assignment $\Theta$, we note that the maximization of $L(\Theta,E,W;\lambda)$ with respect to $\Theta$ under given $E$ and $W$ is equivalent to minimizing
$L(\Theta|E,W;\lambda)=\sum_{j=1}^{p}w_{j}\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{i,i^{\prime}\in c_{k}}\Big{\{}(x_{ij}-E_{ij})-(x_{i^{\prime}j}-E_{i^{\prime}j})\Big{\}}^{2}.$
@@
Given clustering assignment $\Theta$ and variable weight $W$, the error matrix $E$ can be updated as
$\begin{split}{\rm argmin}_{E}\left\{\frac{1}{2}\sum_{j=1}^{p}w_{j}\sum_{k=1}^{K}\sum_{i\in c_{k}}\big{(}x_{ij}-E_{ij}^{\ast}-{\mu}_{kj}^{*}\big{)}^{2}+\sum_{i=1}^{n}P_{1}(\|E_{i\cdot}\|_{2};\lambda_{1})\right\},\end{split}$ (6)
@@
As previously discussed, the variable $E$ plays a crucial role in robustness. If an observation does not fall into any cluster, the presence of $E_{i\cdot}$ will lead to the assumption that this observation pertains to a particular cluster. The penalty $P_{1}(\|E_{i\cdot}\|_{2};\lambda_{1})$ controls the robustness of the model. We also note that under $\lambda_{1}\to\infty$, all $E_{i\cdot}$ equal a $p$-dimensional zero vector. As a result, the loss function (6) will be equivalent to traditional $K$-means clustering.
@@
In the minimization problem (6), as previously mentioned, we employ the group lasso penalty and group SCAD penalty (Tibshirani, 1996; Antoniadis and Fan, 2001). The optimizers are obtained by applying the thresholding function corresponding to the group penalty to $x_{i\cdot}-\mu_{k\cdot}$. These two types of group penalty functions correspond to two distinct categories of thresholding functions: the group lasso penalty function corresponds to the multivariate soft-threshold operator, where $E_{i\cdot}$ can be calculated through
$\begin{split}E_{i\cdot}=(x_{i\cdot}-\mu_{k\cdot})\max\left(0,1-\frac{\lambda_{1}}{\|x_{i\cdot}-\mu_{k\cdot}\|_{2}}\right)\end{split}$ (7)
@@
as proposed by Witten (2013), while the group SCAD penalty function corresponds to the multivariate SCAD-threshold operator introduced by Huang et al. (2012), which $E_{i\cdot}$ can be calculated as:
$\begin{split}E_{i\cdot}=\begin{cases}\mathbf{S}(\mathbf{z};\lambda_{1}),&\text{if }\|\mathbf{z}\|_{2}\leq 2\lambda_{1},\\ \frac{a-1}{a-2}\mathbf{S}(\mathbf{z};\frac{a\lambda_{1}}{a-1}),&\text{if }2\lambda_{1}<\|\mathbf{z}\|_{2}\leq a\lambda_{1},\\ \mathbf{z},&\text{if }\|\mathbf{z}\|_{2}>a\lambda_{1}.\end{cases}\end{split}$ (8)
@@
where $\mathbf{z}=x_{i\cdot}-\mu_{k\cdot}$ and $\mathbf{S}(\mathbf{z};\lambda)$ represents multivariate soft-threshold operator. In this paper, we set $a$ as equal to 3.7, as recommended in Fan and Li (2001).
@@
In addition to the previously mentioned multivariate soft-threshold operator and multivariate SCAD-threshold operator, The application of the aforementioned methods can also be extended to the group MCP penalty associated with the multivariate MCP-threshold operator. However, due to its similarity to the SCAD penalty, we exclusively employed the SCAD penalty. Moreover, we suggest avoiding the application of the group hard penalty as optimizing it is inherently difficult. Furthermore, during our simulation process, we discovered that it is easy to converge to local minima.
@@
As discussed by Witten (2013) and She and Owen (2011), this formulation establishes a connection between the framework of robust M-estimation and the proposed model (6) with clustering center ${\mu}_{kj}^{*}$ for each $k$ held fixed.
@@
As described above, the matrix of errors acquired from the algorithm is weighted based on the current weights. Therefore, it is necessary to restore the weighted error matrix $E$ to its original unweighted state by dividing the current weight before we go to the next step to calculate the new weight for each variable (if the weight of a certain variable is zero, then $E_{\cdot j}$ will divide one instead of zero).
@@
Finally, given $E$ and $\Theta$, the minimization of $L(\Theta,E,W;\lambda)$ with respect to $W$ is equivalent to
@@
$\begin{split}{\rm argmax}_{W\in\mathcal{H}}\left\{\sum_{j=1}^{p}w_{j}Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j})-\sum_{j=1}^{p}(P_{2}(w_{j};\lambda_{2})+\frac{1}{2}w_{j}^{2})\right\},\end{split}$ (9)
@@
where =$\mathcal{H}=\left\{{w_{j}\in\mathbb{R}^{+}:\sqrt{\sum_{j=1}^{p}w^{2}_{j}}=1}\right\}$. We show in the Appendix that the final solution to (9) can be donated as:
$w_{j}=\frac{S(Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j});\lambda_{2})}{\sqrt{\sum_{j=1}^{p}(S(Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j});\lambda_{2}))^{2}}}$

### v2

@@ Abstract
While $K$-means is known to be a standard clustering algorithm, its performance may be compromised due to the presence of outliers and high-dimensional noisy variables. This paper proposes adaptively robust and sparse $K$-means clustering (ARSK) to address these practical limitations of the standard $K$-means algorithm. For robustness, we introduce a redundant error component for each observation, and this additional parameter is penalized using a group sparse penalty. To accommodate the impact of high-dimensional noisy variables, the objective function is modified by incorporating weights and implementing a penalty to control the sparsity of the weight vector. The tuning parameters to control the robustness and sparsity are selected by $\rm Gap$ statistics. Through simulation experiments and real data analysis, we demonstrate the proposed method’s superiority to existing algorithms in identifying clusters without outliers and informative variables simultaneously.
Hao Li1, Shonosuke Sugasawa2 and Shota Katayama2
1Graduate School of Economics, Keio University
2Faculty of Economics, Keio University
@@ Introduction
Identifying clustering structures is recognized as an important task in pursuing potential heterogeneity in datasets. While there are a variety of clustering methods, $K$-means clustering (Forgy, 1965) is the most standard clustering algorithm and is widely used in many scientific applications. However, real-world datasets present significant difficulties, such as the presence of outliers and high-dimensional noisy variables. Regarding the outlier issues, there are some attempts to robustify the standard $K$-means such as trimmed $K$-means (Cuesta-Albertos et al., 1997) and robust $K$-means (Klochkov et al., 2021). On the other hand, high-dimensional noisy characteristics are typically addressed through clustering approaches that focus on sparsity for variable selection. In particular, Witten and Tibshirani (2010) proposed a lasso-type penalty to the variable weights incorporated in the clustering objective function. However, the efficacy of this algorithm might be lost if the dataset contains a significant number of outliers. While several methods address either of the two aspects (the existence of outliers and noisy variables), practically useful methods to address both aspects simultaneously are scarce. One exception is the method proposed by Kondo et al. (2016). However, this approach requires the assumption of a trimming level in advance. In contrast, Brodinová et al. (2019) introduces an approach beginning with eliminating outliers identified by the model. However, this methodology potentially leads to reduced sample size and may result in information loss. Moreover, the excessive variations in gradient magnitudes may destabilize when optimizing the variable weight process unless all outliers are eliminated in advance.
@@
In addition to clustering methods based on objective functions, probabilistic clustering using mixture models is also popular. It is typically done by fitting multivariate Gaussian mixtures. While several approaches have been proposed for stable estimation and clustering under existence of outliers (e.g. Coretto and Hennig, 2016; Fujisawa and Eguchi, 2006; Punzo and McNicholas, 2016; Sugasawa and Kobayashi, 2022; Yang et al., 2012), such approach suffers from a huge number of parameters (especially for modelling covariance matrix) to be estimated when the dimension is large. Therefore, these methods would not be a reasonable solution when our primary focus is on clustering.
@@
To address the difficulty of a series of sparse trimming-cluster and probabilistic-cluster approaches, we propose a new approach based on regularization techniques for sparsity as well as robustness, called adaptively robust and sparse $K$-means (ARSK). We introduce an error component for each observation as an additional parameter to absorb the undesirable effects caused by outliers, making the clustering algorithm robust. We employ a group lasso penalty and a group smoothly clipped absolute deviation (SCAD) penalty (Wang et al., 2007) to estimate these parameters during clustering steps. Meanwhile, to reduce the data dimension, we also introduce a SCAD penalty (Fan and Li, 2001) to weights for each variable to eliminate noise variables irrelevant to clustering. Based on our analysis of real-world datasets, in the ultra-high dimensional setting (i.e., the dataset dimension exceeding 700), the result procured through the implementation of the lasso penalty proposed by Witten and Tibshirani (2010) frequently poor performance. Therefore, we have developed a methodology based on the SCAD penalty. The strong sparsity property of the SCAD penalty effectively improves the selection of relevant variables in the ultra-high dimensional scenario. We then develop an efficient computation algorithm to minimize the proposed objective function. While the proposed method has two tuning parameters controlling robustness (the number of outliers to be detected) and sparsity (the level of the sparsity of the variable weight), we adopt a modified version of the $\rm Gap$ statistics (Tibshirani et al., 2001) to determine the optimal values of them. Hence, the level of robustness and sparsity is adaptively determined in a fully data-dependent manner.
@@
The paper is organized as follows. Section 2 reviews the $K$-means algorithm and its extended model introduces the proposed method ARSK and presents its details and pseudocode. We also present the selection approach of tuning parameters and its pseudocode. while Section 3 presents the tuning parameter selection result. Both two algorithms are applied to various scenarios of artificial datasets and several real-world datasets. This paper ends in Section 4 with concluding remarks. R code implementing the proposed method is available at the GitHub repository (https://github.com/lee1995hao/ARSK).
@@
Suppose that we observe a $p$-dimensional vector, $\bm{X}_{i,:}=(X_{i1},\ldots,X_{ip})$, for $i=1,\ldots,n$, and that we are interested in clustering $\bm{X}_{i,:}$. Given the number of clusters, $K$, the conventional $K$-means algorithm is to find the optimal clustering by minimizing the following objective function:
$\min_{c_{1},...,c_{K}}\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{{i},{i^{\prime}}\in c_{k}}\|\bm{X}_{i,:}-\bm{X}_{i^{\prime},:}\|^{2},$ (1)
where $\|\bm{X}_{i,:}-\bm{X}_{i^{\prime},:}\|^{2}$ is the squared value of $L_{2}$-distance between two observed vectors. Here $c_{k}$ is a set of indices representing a cluster of observations and $n_{k}$ is the size of $c_{k}$, such that $\cup_{k=1}^{K}c_{k}=\{1,\ldots,n\}$ and $\sum_{k=1}^{K}n_{k}=n$, where $c_{k}\cap c_{k^{\prime}}=\phi$ for $k\neq k^{\prime}$. The objective function (1) can be easily optimized by iteratively updating cluster means and assignment (e.g. Lloyd, 1982).
As demonstrated in Witten and Tibshirani (2010), the objective function (1) can be reformulated in terms of the between-cluster sum of squares, given by $\max_{c_{1},..,c_{k}}\sum_{j=1}^{p}Q_{j}(\bm{X}_{:,j};\Theta)$, where
$Q_{j}(\bm{X}_{:,j};\Theta)=\frac{1}{n}\sum_{i=1}^{n}\sum_{i^{\prime}=1}^{n}(X_{ij}-X_{i^{\prime}j})^{2}-\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{i,i^{\prime}\in c_{k}}(X_{ij}-X_{i^{\prime}j})^{2},$ (2)
with $\bm{X}_{:,j}=(X_{1j},\ldots,X_{nj})$ and $\Theta$ being a collection of $c_{1},\ldots,c_{K}$. Witten and Tibshirani (2010) employed the modified version of the objective function (2), given by $\sum_{j=1}^{p}\omega_{j}Q_{j}(\bm{X}_{:,j};\Theta)$ with an appropriate constrained on $\omega_{j}$, for conducting variable selection and clustering simultaneously.
@@
The objective function (2) is sensitive to outliers since the approach relies on the $L_{2}$ distance between each observation. As a result, even a small quantity of deviating observations can affect the $K$-means algorithm. In order to robustify the $K$-means approach, we modify the function $Q_{j}(\bm{X}_{:,j};\Theta)$ as follows:
$\centering\begin{split}Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E})&=\frac{1}{n}\sum_{i=1}^{n}\sum_{i^{\prime}=1}^{n}\{(X_{ij}-E_{ij})-(X_{i^{\prime}j}-E_{i^{\prime}j})\}^{2}\\ &\ \ \ -\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{i,i^{\prime}\in c_{k}}\{(X_{ij}-E_{ij})-(X_{i^{\prime}j}-E_{i^{\prime}j})\}^{2},\end{split}\@add@centering$ (3)
where $\bm{E}$ is an $n\times p$ error matrix, which is a collection of $E_{ij}$ and $E_{ij}$ is an additional location parameter for each $X_{ij}$ such that $E_{ij}$ will be non-zero values for outlying observations $X_{ij}$ to prevent an undesirable effects from outliers. While the number of elements of $\bm{E}$ is the same as the number of observations, it can be assumed that $\bm{E}$ is sparse in the sense that most elements of $\bm{E}$ are 0. That is, most elements are not outliers. To stably estimate $\bm{E}$, we consider penalized estimation for $E_{ij}$ when optimizing the function (3). Such approaches are used in the robust fitting of regression models (e.g. She and Owen, 2011; Katayama and Fujisawa, 2017).
@@
$L(\Theta,\bm{E},\bm{w};\lambda)=\sum_{j=1}^{p}w_{j}Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j})-\sum_{i=1}^{n}P_{1}(\|\bm{E}_{i,:}\|_{2};\lambda_{1})-\sum_{j=1}^{p}(P_{2}(w_{j};\lambda_{2})+\frac{1}{2}w_{j}^{2}),$ (4)
where $P_{1}$ and $P_{2}$ are group penalty functions and penalty functions for sparse (shrinkage) estimation of $\bm{E}_{i,:}$ and $w_{j}$, respectively, and $\lambda=(\lambda_{1},\lambda_{2})$ is a set of tuning parameters. To find a unique solution for the weight coefficients $w_{j}$, we add a quadratic term $\frac{1}{2}\sum_{j=1}^{p}w_{j}^{2}$ to the objective function. Then, the robust and sparse $K$-means clustering can be defined as $\widehat{\Theta}$ such that
$\begin{split}(\widehat{\Theta},\widehat{\bm{E}},\widehat{\bm{w}})={\rm argmax}_{\Theta,\bm{E},\bm{w}\in\mathcal{H}}L(\Theta,\bm{E},\bm{w};\lambda),\\ \mathcal{H}=\left\{{w_{j}\in\mathbb{R}^{+}:\sqrt{\sum_{j=1}^{p}w^{2}_{j}}=1}\right\}.\end{split}$
In what follows, we employ two specific forms of group penalty functions $P_{1}(\|\bm{E}_{i,:}\|_{2};\lambda_{1})$: one type is the convex group penalty function, which is the group lasso penalty as $\lambda_{1}\|\bm{E}_{i,:}\|_{2}$ (Yuan and Lin, 2006), and the other type is the nonconvex group penalty function, which is the group SCAD penalty (Huang et al., 2012) as
$P_{\lambda_{1}}^{SCAD}\left(\|\bm{E}_{i,:}\|_{2}\right)=\left\{\begin{array}[]{ll}\lambda_{1}\|\bm{E}_{i,:}\|_{2}&\text{ if }\|\bm{E}_{i,:}\|_{2}\leq\lambda_{1};\\ -(\|\bm{E}_{i,:}\|_{2}^{2}-2a\lambda_{1}\|\bm{E}_{i,:}\|_{2}+\lambda_{1}^{2})/2(a-1)&\text{ if }\lambda_{1}<\|\bm{E}_{i,:}\|_{2}\leq a\lambda_{1};\\ (a+1)\lambda_{1}^{2}/2&\text{ if }\|\bm{E}_{i,:}\|_{2}>a\lambda_{1}.\end{array}\right.$
@@
$P_{\lambda_{2}}^{SCAD}\left(w_{j}\right)=\left\{\begin{array}[]{ll}\lambda_{2}\left|w_{j}\right|&\text{ if }\left|w_{j}\right|\leq\lambda_{2};\\ -(\left|w_{j}\right|^{2}-2a\lambda_{2}\left|w_{j}\right|+\lambda_{2}^{2})/2(a-1)&\text{ if }\lambda_{2}<\left|w_{j}\right|\leq a\lambda_{2};\\ (a+1)\lambda_{2}^{2}/2&\text{ if }\left|w_{j}\right|>a\lambda_{2}.\end{array}\right.$
When applying the $L_{1}$ or SCAD penalties, we seek the optimal weight vector of $\bm{w}=(w_{1},\ldots,w_{p})$, which is normalized to have a unit $L_{2}$-norm, i.e., $\sqrt{\sum_{j=1}^{p}w^{2}_{j}}=1$. This approach helps to promote a more balanced solution (Zou and Hastie, 2010).
@@
The block-coordinated decent method can optimise the proposed objective function (4) for updating parameters. Regarding the update of the clustering assignment $\Theta$, we note that the maximization of $L(\Theta,\bm{E},\bm{w};\lambda)$ with respect to $\Theta$ given $\bm{E}$ and $\bm{w}$ is equivalent to minimizing
$L(\Theta|\bm{E},\bm{w};\lambda)=\sum_{j=1}^{p}w_{j}\sum_{k=1}^{K}\frac{1}{n_{k}}\sum_{i,i^{\prime}\in c_{k}}\Big{\{}(X_{ij}-E_{ij})-(X_{i^{\prime}j}-E_{i^{\prime}j})\Big{\}}^{2}.$
@@
Given clustering assignment $\Theta$ and variable weight $\bm{w}$, the error matrix $\bm{E}$ can be updated as
$\begin{split}{\rm argmin}_{E}\left\{\frac{1}{2}\sum_{j=1}^{p}w_{j}\sum_{k=1}^{K}\sum_{i\in c_{k}}\big{(}X_{ij}-E_{ij}^{\ast}-{\mu}_{kj}^{*}\big{)}^{2}+\sum_{i=1}^{n}P_{1}(\|\bm{E}_{i,:}\|_{2};\lambda_{1})\right\},\end{split}$ (6)
@@
As previously discussed, the variable $\bm{E}$ plays a crucial role in robustness. If an observation does not fall into any cluster, vector $\bm{E}_{i,:}$ elements will exhibit large values. The penalty $P_{1}(\|\bm{E}_{i,:}\|_{2};\lambda_{1})$ controls the robustness of the model. We also note that under $\lambda_{1}\to\infty$, all $\bm{E}_{i,:}$ equal a $p$-dimensional zero vector. As a result, the objective function (6) will be equivalent to traditional $K$-means clustering.
@@
In the minimization problem (6), as previously mentioned, we employ the group lasso penalty and group SCAD penalty (Tibshirani, 1996; Antoniadis and Fan, 2001). The optimizers are obtained by applying the thresholding function corresponding to the group penalty to $\bm{X}_{i,:}-\bm{\mu}_{k,:}$. These two types of group penalty functions correspond to two distinct categories of thresholding functions: the group lasso penalty function corresponds to the multivariate soft-threshold operator, where $\bm{E}_{i,:}$ can be calculated through
$\begin{split}\bm{E}_{i,:}=(\bm{X}_{i,:}-\bm{\mu}_{k,:})\max\left(0,1-\frac{\lambda_{1}}{\|\bm{X}_{i,:}-\bm{\mu}_{k,:}\|_{2}}\right)\end{split}$ (7)
@@
as proposed by Witten (2013), while the group SCAD penalty function corresponds to the multivariate SCAD-threshold operator introduced by Huang et al. (2012), which $\bm{E}_{i,:}$ can be calculated as:
$\begin{split}\bm{E}_{i,:}=\begin{cases}\bm{S}(\bm{z}_{i};\lambda_{1}),&\text{if }\|\bm{z}_{i}\|_{2}\leq 2\lambda_{1},\\ (a-2)^{-1}(a-1)\bm{S}(\bm{z}_{i};(a-1)^{-1}a\lambda_{1}),&\text{if }2\lambda_{1}<\|\bm{z}_{i}\|_{2}\leq a\lambda_{1},\\ \bm{z}_{i},&\text{if }\|\bm{z}_{i}\|_{2}>a\lambda_{1}.\end{cases}\end{split}$ (8)
@@
where $\bm{z}_{i}=\bm{X}_{i,:}-\bm{\mu}_{k,:}$ and $\bm{S}(\bm{z}_{i};\lambda)$ represents multivariate soft-threshold operator. In this paper, we set $a$ as equal to 3.7, as recommended in Fan and Li (2001).
@@
In addition to the previously mentioned multivariate soft-threshold operator and multivariate SCAD-threshold operator, The application of the aforementioned methods can also be extended to the group minimax concave penalty (Huang and Zhang, 2010) associated with the multivariate minimax concave penalty-threshold operator. However, due to its similarity to the SCAD penalty, we exclusively employed the SCAD penalty. Moreover, we suggest avoiding applying the group hard penalty (Antoniadis, 1996), as optimizing it is inherently difficult. Furthermore, during our simulation process, we discovered that it is easy to converge to local minima.
@@
As discussed by Witten (2013) and She and Owen (2011), this formulation establishes a connection between the framework of robust M-estimation and the proposed model (6) with clustering centre ${\mu}_{kj}^{*}$ for each $k$ held fixed.
@@
As described above, the matrix of errors acquired from the algorithm is weighted based on the current weights. Therefore, it is necessary to restore the weighted error matrix $\bm{E}$ to its original unweighted state by dividing the current weight before we go to the next step to calculate the new weight for each variable (if the weight of a certain variable is zero, then $\bm{E}_{:,j}$ will divide one instead of zero).
@@
Finally, given $\bm{E}$ and $\Theta$, the minimization of $L(\Theta,\bm{E},\bm{w};\lambda)$ with respect to $\bm{w}$ is equivalent to
@@
$\begin{split}{\rm argmax}_{\bm{w}\in\mathcal{H}}\left\{\sum_{j=1}^{p}w_{j}Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j})-\sum_{j=1}^{p}\left(P_{2}(w_{j};\lambda_{2})+\frac{1}{2}w_{j}^{2}\right)\right\},\end{split}$ (9)
@@
where =$\mathcal{H}=\left\{{w_{j}\in\mathbb{R}^{+}:\sqrt{\sum_{j=1}^{p}w^{2}_{j}}=1}\right\}$. We show in the Appendix that the final solution to (9) can be denoted as
$w_{j}=\frac{S(Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j});\lambda_{2})}{\sqrt{\sum_{j=1}^{p}(S(Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j});\lambda_{2}))^{2}}}.$